OpenStudy (lovelyharmonics):

Evaluate the expression: v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5>

OpenStudy (luigi0210):

Haven't done vectors in a while.. >.<

OpenStudy (lovelyharmonics):

noooo :c

OpenStudy (praxer):

v=2i-8j-8k w=-2i+6j-5k since v.w=vwcostheta. Therefore for cos theta = 90 it will be zero. So multiply the coefficients of the corresponding directions. we have v.w = -4-48+40=-4-8=-12.. So v dot w = -12...

OpenStudy (praxer):

## sorry for theta = 90 cos theta is 0

OpenStudy (praxer):

theta is the angle between the direction axis...